wat is the solution to this equation?
n + 18 + 2n = -12
a.) n = -10
b.) n = -2
c.) n = 1/2
d.) n = 6
e.) n = 9

Answer by longjonsilver(2297)

with any equation, you need to collect like terms together and follow some basic rules like "whatever you do on one side of the = you have to do on the other side too". That is about it really.
so, n + 18 + 2n = -12
collecting all the n terms together gives 3n + 18 = -12
now we want to remove the +18 from the lefthand side, so to do this we take away 18. Here the rule above kicks in...we have done something to the lefthand side, so we must do the same to the righthand side: In full it is: 3n+18-18=-12-18. But we never write this. We just write the following line
3n = -30
So this says 3 "n's" are 30, so what is 1 n?. Answer?
We have 3 of these "n's" and we want to make it just 1 of them whilst leaving the n alone, so we divide the 3 by 3 to give 1. And we do the same to other side...-30 divided by 3 is -10....
n = -10
Hope this is somewhat clearer?
cheers
Jon